8. An aluminum wire carrying a current of 5 A has a cross sectional area of 4 x 106 m2. Find the drift speed of the electrons in the wire. The density of aluminum is 2.7 g/cm3. (Assume three electrons are supplied by each atom). a) Recall the mass flow rate equation from semester 1. Rms Ro v b) Write the equivalent current relation for charges in a wire, I- c) What is the electrical analogy with fluid density, p? d) Given that the molar mass of aluminum is M 27 g/mol and the density of aluminum is drift mol/cm3 ρ-27 g/cm3, determine the number of mol cm3 for aluminum. ρ/M Convert into atoms/m3, N/V-p M (mol/cm3) x(10cm3/m3) (6.02E23 atoms mol) = e) atoms/m esnr edoom criutes 3 conducting electrons. n = 3(electrons/atom) ×N/V (atoms/m3) g) Get the charge density, nq h) Solve the equation in b) for the drift velocity, Vdrit i) Plug in the numbers to get electrons/m3 C/m Vdriftー m/s j) How long would it take an electron to drift 1 meter through this wire? t- sec all of 8 please.
Answer #1

a)Mass flow rate is  R_{mass\, flow}=\rho *vA  

Where \rho is density

b) Current I=nq\, v_{drift}A

c) Electrical analogy with fluid density \rho is charge density  nq

d) \rho/M=\frac{2.7}{27}=0.1\,mol/cm^3

e) N/V-P/M*106 6.02102010 6.02 102 .0102 atom/m

f) n=3*N/V=3*6.02*10^{28}=1.806*10^{29}\,electrons/m^3

g)Charge density nq=1.806*10^{29}*1.6*10^{-19}=2.9*10^{10}\,C/m^3

h)I=nq\, v_{drift}A

v_{drift}=\frac{I}{nq A}

i)

-4.33 10 m/s Udrift nq nqA 2.91014 10-6

j) Time taken Udift 4.33 10-5 2.31 * 104s